# Differentiability of solutions for a degenerate fully nonlinear free transmission problem

### v1

@@ 4.1 Proof of Theorem 2
$\frac{7}{10}\sum\limits_{i=1}^{\infty}\sigma_{2}^{-1}(\theta^{k})\leq\sum\limits_{i=1}^{\infty}\frac{\sigma_{2}^{-1}(\theta^{k})}{c_{k}}\leq\sum\limits_{i=1}^{\infty}\sigma_{2}^{-1}(\theta^{k}).$
@@
$\begin{split}\sigma_{i}^{0}(t)&:=\sigma_{i}(t),\\ \sigma_{i}^{1}(t)&:=\frac{\mu_{1}}{r}\sigma_{i}(\mu_{1}t),\\ \sigma_{i}^{2}(t)&:=\frac{\mu_{1}\mu_{2}}{r^{2}}\sigma_{i}(\mu_{1}\mu_{2}t),\\ &\vdots\\ \sigma_{i}^{k}(t)&:=\frac{\prod\limits_{j=1}^{k}\mu_{j}}{r^{k}}\sigma_{i}(\prod\limits_{j=1}^{k}\mu_{j}t),\end{split}$
with $\mu_{1}>r$ as defined and $(\mu_{k})_{k\in\mathbb{N}}$ determined as follows. If
$\frac{\mu_{k}\prod\limits_{j=1}^{k-1}\mu_{j}}{r^{k}}\sigma_{2}\left(\mu_{k}\prod\limits_{j=1}^{k-1}\mu_{j}c_{k}\right)\geq 1,$
then $\mu_{k}=\mu_{k-1}$. Otherwise $\mu_{k}<1$ is chosen to ensure
$\frac{\prod\limits_{j=1}^{k}\mu_{j}}{r^{k}}\sigma_{2}\left(\prod\limits_{j=1}^{k}\mu_{j}c_{k}\right)=1.$
We notice Assumption 3 ensures that such a choice for the sequence $(\mu_{k})_{k\in\mathbb{N}}$ implies
$\frac{\prod\limits_{j=1}^{k}\mu_{j}}{r^{k}}\sigma_{1}\left(\prod\limits_{j=1}^{k}\mu_{j}c_{k}\right)\geq 1.$
Once these ingredients are available, we combine them with Proposition 9 to produce a sequence of affine functions whose difference with respect to $u$ grows in a controlled fashion.
@@
Let $u\in C(B_{1})$ be a normalized viscosity solution to (19)-(20). Suppose Assumptions 1-4 hold true. There exists $\varepsilon>0$ such that, if $\left\|f\right\|_{L^{\infty}(B_{1})}<\varepsilon$, one finds $a\in\mathbb{R}$ and ${\rm b}\in\mathbb{R}^{d}$ satisfying $|a|+\left\|b\right\|\leq C$, for some universal constant $C>0$. In addition,
@@
Let $u\in C(B_{1})$ be a normalized viscosity solution of (19)-(20). Assume that Assumptions 1-4 hold and that $\left\|f\right\|_{L^{\infty}(B_{1})}\leq\varepsilon$, where $\varepsilon$ is the same as in Proposition 10. Then for every $n\in\mathbb{N}$, there are affine functions $(\phi_{n})_{n\in\mathbb{N}}$ of the form
$\phi_{n}(x):=a_{n}+{\rm b}_{n}\cdot x$ (34)
@@
$\sup_{x\in B_{r_{n}}}|u(x)-\phi_{n}(x)|\leq\left(\prod\limits_{i=1}^{n}\mu_{i}\right)r_{n},$ (35)
$|a_{n+1}-a_{n}|\leq C\left(\prod\limits_{i=1}^{n}\mu_{i}\right)r_{n},$ (36)
@@
The proof follows from an induction argument. The case $n=1$ is covered by taking $\phi_{1}:=a+{\rm b}\cdot x$, where $a\in\mathbb{R}$ and ${\rm b}\in\mathbb{R}^{d}$ are as in Proposition 10. It accounts for the base case.
@@
$u_{k}(x):=\frac{u_{k-1}(rx)-\ell_{k-1}(rx)}{\mu_{k}\cdot r},$ (38)
where $r<\mu_{1}\leq\mu_{2}\ldots\leq\mu_{k}$ have been chosen earlier. Note that $u_{k}$ solves
@@
$F_{i}^{k}(p,M):=\sigma_{i}^{k}\left(\left|p+\frac{D\phi_{k-1}}{\mu_{k-1}}\right|\right)F_{k}(M),$
@@
$\sigma_{i}^{k}(t):=\frac{\prod\limits_{j=1}^{k}\mu_{j}}{r^{k}}\sigma_{i}\left(\prod\limits_{j=1}^{k}\mu_{j}t\right),$
$F_{k}(M):=r^{k}\left(\prod\limits_{j=1}^{k}\mu_{j}\right)^{-1}F\left(\left(\prod\limits_{j=1}^{k}\mu_{j}\right)\left(r^{k}\right)^{-1}M\right),$
@@
$f_{k}(x):=f(r^{k}x).$
The choice of $(\mu_{j})_{j=1}^{k}$ and the construction of the degeneracies $\sigma_{i}^{k}$ allows us to evoke Proposition 10 and obtain an affine function $\phi$ satisfying
@@
$\sup_{x\in B_{r^{k}}}|u(x)-\phi_{k}(x)|\leq\left(\prod\limits_{i=1}^{k}\mu_{i}\right)r^{k},$
@@
$\phi_{k}(x):=\phi_{1}(x)+\sum\limits_{i=2}^{k-1}\phi_{i}(r^{-1}x)\prod\limits_{i=1}^{k-1}(\mu_{i})r^{i}=a_{k}+{\rm b}_{k}\cdot x.$
@@
$|a_{k}-a_{k-1}|\leq C\left(\prod\limits_{i=1}^{k-1}\mu_{i}\right)r^{k-1},$
@@
$|{\rm b}_{k}-{\rm b}_{k-1}|\leq C\left(\prod\limits_{i=1}^{k-1}\mu_{i}\right),$
@@
We start by noticing two possibilities concerning the sequence $(\mu_{n})_{n\in\mathbb{N}}$. Either the sequence repeats after some index $N\geq 2$ or we have $\mu_{n}<\mu_{n+1}$ for infinitely many indices $n\in\mathbb{N}$.
@@
$\frac{\prod\limits_{i=1}^{n+1}\mu_{i}}{r^{n+1}}\sigma_{2}\left(\left[\prod\limits_{i=1}^{n+1}\mu_{i}\right]c_{n+1}\right)=1.$
Here, the definition of $\mu_{n}$ implies
$\frac{\prod\limits_{i=1}^{n}\mu_{i}}{r^{n}}\sigma_{2}\left(\prod\limits_{i=1}^{n}\mu_{i}\,c_{n}\right)=1.$
@@
$\prod\limits_{i=1}^{n}\mu_{i}=\frac{1}{c_{n}}\sigma_{2}^{-1}\left(\frac{r^{n}}{\prod\limits_{i=1}^{n}\mu_{i}}\right)\leq\frac{\sigma_{2}^{-1}(\theta_{n})}{c_{n}}.$ (39)
@@
$(\tau_{n})_{n\in\mathbb{N}}:=\left(\prod\limits_{i=1}^{n}\mu_{i}\right)_{n\in\mathbb{N}}.$
Because of (39), we conclude that $(\tau_{n})_{n\in\mathbb{N}}\in\ell^{1}$, and its norm is bounded by $\sum\limits_{i=1}^{\infty}\sigma_{2}^{-1}(\theta_{i})$. The latter is finite due to the summable characterisation of Dini continuity and Assumption 4.
As a consequence, we infer that
$\lim_{n\to\infty}\left(\prod\limits_{i=1}^{n}\mu_{i}\right)=0;$
@@
$|a_{\infty}-a_{n}|\leq C\left(\sum\limits_{i=n}^{\infty}\tau_{i}\right)r^{n}\qquad\mbox{and}\qquad|{\rm b}_{\infty}-{\rm b}_{n}|\leq C\left(\sum\limits_{i=n}^{\infty}\tau_{i}\right).$ (40)
Set $\varphi(x):=a_{\infty}+{\rm b}_{\infty}\cdot x$ and fix $0<\rho\ll 1$. Let $n\in\mathbb{N}$ be such that $r^{n}<\rho\leq r^{n+1}$. Combine Proposition 11 with the inequalities in (40) to obtain
$\begin{split}\sup_{x\in B_{\rho}}\left|u(x)-\varphi(x)\right|&\leq\sup_{x\in B_{r^{n}}}\left|u(x)-\phi_{n}(x)\right|+\sup_{x\in B_{r^{n}}}\left|\varphi(x)-\phi_{n}(x)\right|\\ &\leq\frac{1}{r}C\left(\tau_{n}+\sum_{i=n}^{\infty}\tau_{i}\right)\rho\\ &\leq C\left(\sum_{i=n}^{\infty}\tau_{i}\right)\rho.\end{split}$ (41)
@@
$n(\rho):=\left\lfloor 1/\rho\right\rfloor.$
@@
$\sigma(t):=\sum_{i=\left\lfloor 1/t\right\rfloor}^{\infty}\tau_{i}$
@@
Acknowledgements. EP is supported by the Centre for Mathematics of the University of Coimbra (CMUC, funded by the Portuguese Government through FCT/MCTES, DOI 10.54499/UIDB/00324/2020). DS is supported by ISF Grant No. 296/20.

### v2

@@ 3.1. Proof of Theorem 1
$\frac{7}{10}\sum\limits_{i=1}^{\infty}a_{k}\leq\sum\limits_{i=1}^{\infty}\frac{a_{k}}{c_{k}}\leq\sum\limits_{i=1}^{\infty}a_{k}<\infty.$
@@
$\begin{split}\sigma_{i}^{0}(t)&\coloneqq\sigma_{i}(t),\\ \sigma_{i}^{1}(t)&\coloneqq\frac{\mu_{1}}{r}\sigma_{i}(\mu_{1}t),\\ \sigma_{i}^{2}(t)&\coloneqq\frac{\mu_{1}\mu_{2}^{*}}{r^{2}}\sigma_{i}(\mu_{1}\mu_{2}^{*}t),\\ &\vdots\\ \sigma_{i}^{k}(t)&\coloneqq\frac{\mu_{1}\prod\limits_{j=2}^{k}\mu_{j}^{*}}{r^{k}}\sigma_{i}\left(\mu_{1}\prod\limits_{j=2}^{k}\mu_{j}^{*}t\right),\end{split}$
with $\mu_{1}>r$ as defined and $(\mu_{k}^{*})_{k\in\mathbb{N}}$ determined as follows. Set
$\mu_{k}^{*}\coloneqq\max\left(\mu_{k}^{1},\mu^{2}_{k}\right).$
If
$\frac{\mu_{k-1}^{*}\prod\limits_{j=1}^{k-1}\mu^{*}_{j}}{r^{k}}\sigma_{i}\left(\mu_{k-1}^{*}\prod\limits_{j=1}^{k-1}\mu^{*}_{j}c_{k}\right)\geq 1,$
then $\mu_{k}^{i}=\mu_{k-1}^{*}$. Otherwise $\mu_{k}^{i}<1$ is chosen to ensure
$\frac{\mu_{k}^{i}\prod\limits_{j=1}^{k-1}\mu^{*}_{j}}{r^{k}}\sigma_{i}\left(\mu_{k}^{i}\prod\limits_{j=1}^{k-1}\mu_{j}c_{k}\right)=1.$
As before, $\mu_{k}^{*}\coloneqq\max\left(\mu_{k}^{1},\mu^{2}_{k}\right)$. Once these ingredients are available, we combine them with Proposition 6 to produce a sequence of affine functions whose difference with respect to $u$ grows in a controlled fashion.
@@
Let $u\in C(B_{1})$ be a normalized viscosity solution to (7)-(8). Suppose Assumptions 1 and 2 hold true. There exists $\varepsilon>0$ such that, if $\left\|f\right\|_{L^{\infty}(B_{1})}<\varepsilon$, one finds $a\in\mathbb{R}$ and ${\rm b}\in\mathbb{R}^{d}$ satisfying $|a|+\left\|b\right\|\leq C$, for some universal constant $C>0$. In addition,
@@
Let $u\in C(B_{1})$ be a normalized viscosity solution of (7)-(8). Assume that Assumptions 1 and 2 hold and that $\left\|f\right\|_{L^{\infty}(B_{1})}\leq\varepsilon$, where $\varepsilon$ is the same as in Proposition 7. Then for every $n\in\mathbb{N}$, there are affine functions $(\phi_{n})_{n\in\mathbb{N}}$ of the form
$\phi_{n}(x)\coloneqq a_{n}+{\rm b}_{n}\cdot x$
@@
$\sup_{x\in B_{r_{n}}}|u(x)-\phi_{n}(x)|\leq\left(\prod\limits_{j=1}^{n}\mu^{*}_{j}\right)r_{n},$
$|a_{n+1}-a_{n}|\leq C\left(\prod\limits_{j=1}^{n}\mu_{j}^{*}\right)r_{n},$
@@
The proof follows from an induction argument. The case $n=1$ is covered by taking $\phi_{1}\coloneqq a+{\rm b}\cdot x$, where $a\in\mathbb{R}$ and ${\rm b}\in\mathbb{R}^{d}$ are as in Proposition 7. It accounts for the base case.
@@
(20) $u_{k}(x)\coloneqq\frac{u_{k-1}(rx)-\ell_{k-1}(rx)}{\mu^{*}_{k}\cdot r},$
where $r<\mu_{1}\leq\mu_{2}^{*}\ldots\leq\mu_{k}^{*}$ have been chosen earlier. Note that $u_{k}$ solves
@@
$F_{i}^{k}(p,M)\coloneqq\sigma_{i}^{k}\left(\left|p+\frac{D\phi_{k-1}}{\mu^{*}_{k-1}}\right|\right)F_{k}(M),$
@@
$\sigma_{i}^{k}(t)\coloneqq\frac{\prod\limits_{j=1}^{k}\mu^{*}_{j}}{r^{k}}\sigma_{i}\left(\prod\limits_{j=1}^{k}\mu^{*}_{j}t\right),$
$F_{k}(M)\coloneqq r^{k}\left(\prod\limits_{j=1}^{k}\mu^{*}_{j}\right)^{-1}F\left(\left(\prod\limits_{j=1}^{k}\mu^{*}_{j}\right)\left(r^{k}\right)^{-1}M\right),$
@@
$f_{k}(x)\coloneqq f(r^{k}x).$
The choice of $(\mu^{*}_{j})_{j=1}^{k}$ and the construction of the degeneracies $\sigma_{i}^{k}$ allows us to evoke Proposition 7 and obtain an affine function $\phi$ satisfying
@@
$\sup_{x\in B_{r^{k}}}|u(x)-\phi_{k}(x)|\leq\left(\prod\limits_{j=1}^{k}\mu^{*}_{j}\right)r^{k},$
@@
$\phi_{k}(x)\coloneqq\phi_{1}(x)+\sum\limits_{i=2}^{k-1}\phi_{i}(r^{-1}x)\prod\limits_{i=1}^{k-1}\mu_{j}^{*}r^{i}=a_{k}+{\rm b}_{k}\cdot x.$
@@
$|a_{k}-a_{k-1}|\leq C\left(\prod\limits_{j=1}^{k-1}\mu^{*}_{j}\right)r^{k-1},$
@@
$|{\rm b}_{k}-{\rm b}_{k-1}|\leq C\left(\prod\limits_{j=1}^{k-1}\mu^{*}_{j}\right),$
@@
We start by noticing two possibilities concerning the sequence $(\mu^{*}_{n})_{n\in\mathbb{N}}$. Either the sequence repeats after some index $N\geq 2$ or we have $\mu_{n}^{*}<\mu^{*}_{n+1}$ for infinitely many indices $n\in\mathbb{N}$.
@@
$\frac{\prod\limits_{j=1}^{n+1}\mu_{j}^{*}}{r^{n+1}}\sigma_{i}\left(\left[\prod\limits_{j=1}^{n+1}\mu^{*}_{j}\right]c_{n+1}\right)=1.$
Here, the definition of $\mu_{n}^{*}$ implies
$\frac{\prod\limits_{j=1}^{n}\mu^{*}_{j}}{r^{n}}\sigma_{i}\left(\prod\limits_{j=1}^{n}\mu^{*}_{j}\,c_{n}\right)=1.$
@@
(21) $\prod\limits_{j=1}^{n}\mu^{*}_{j}=\frac{1}{c_{n}}\sigma_{j}^{-1}\left(\frac{r^{n}}{\prod\limits_{j=1}^{n}\mu^{*}_{j}}\right)\leq\frac{\sigma_{i}^{-1}(\theta_{n})}{c_{n}}.$
@@
$(\tau_{n})_{n\in\mathbb{N}}\coloneqq\left(\prod\limits_{j=1}^{n}\mu^{*}_{j}\right)_{n\in\mathbb{N}}.$
Notice that
$\begin{split}\prod_{j=1}^{m+1}\mu_{j}^{*}&=\frac{1}{c_{m+1}}\sigma_{i}^{-1}\left(\frac{r^{m+1}}{\prod_{i=1}^{m+1}\mu_{i}^{*}}\right)\leq\frac{\sigma_{i}^{-1}\left(\theta^{m+1}\right)}{c_{m+1}}\\ &\leq\frac{\max\left(\sigma_{1}^{-1}\left(\theta^{m+1}\right),\sigma_{2}^{-1}\left(\theta^{m+1}\right)\right)}{c_{m+1}}.\end{split}$
Because of (21), we conclude that $(\tau_{n})_{n\in\mathbb{N}}\in\ell^{1}$, and its norm is bounded by
$\sum_{n=1}^{\infty}\max\left(\frac{\sigma_{1}^{-1}(\theta_{n})}{c_{n}},\frac{\sigma_{2}^{-1}(\theta_{n})}{c_{n}}\right)$
The latter is finite due to the summable characterisation of Dini continuity and Assumption 2. As a consequence, we infer that
$\lim_{n\to\infty}\left(\prod\limits_{k=1}^{n}\mu_{k}^{*}\right)=0;$
@@
(22) $|a_{\infty}-a_{n}|\leq C\left(\sum\limits_{i=n}^{\infty}\tau_{i}\right)r^{n}\qquad\mbox{and}\qquad|{\rm b}_{\infty}-{\rm b}_{n}|\leq C\left(\sum\limits_{i=n}^{\infty}\tau_{i}\right).$
Set $\varphi(x)\coloneqq a_{\infty}+{\rm b}_{\infty}\cdot x$ and fix $0<\rho\ll 1$. Let $n\in\mathbb{N}$ be such that $r^{n}<\rho\leq r^{n+1}$. Combine Proposition 8 with the inequalities in (22) to obtain
(23) $\begin{split}\sup_{x\in B_{\rho}}\left|u(x)-\varphi(x)\right|&\leq\sup_{x\in B_{r^{n}}}\left|u(x)-\phi_{n}(x)\right|+\sup_{x\in B_{r^{n}}}\left|\varphi(x)-\phi_{n}(x)\right|\\ &\leq\frac{1}{r}C\left(\tau_{n}+\sum_{i=n}^{\infty}\tau_{i}\right)\rho\\ &\leq C\left(\sum_{i=n}^{\infty}\tau_{i}\right)\rho.\end{split}$
@@
$n(\rho)\coloneqq\left\lfloor 1/\rho\right\rfloor.$
@@
$\sigma(t)\coloneqq\sum_{i=\left\lfloor 1/t\right\rfloor}^{\infty}\tau_{i}$
@@
Acknowledgements. This publication is based upon work supported by King Abdullah University of Science and Technology Research Funding (KRF) under Award No. ORFS-2024-CRG12-6430.3. EP is supported by the Centre for Mathematics of the University of Coimbra (CMUC, funded by the Portuguese Government through FCT/MCTES, DOI 10.54499/UIDB/00324/2020). DS is supported by the National Science Centre (NCN) Grant Sonata Bis 2019/34/E/ST1/00120.
Declarations
Data availability statement: All data needed are contained in the manuscript.
Funding and/or Conflicts of interests/Competing interests: The authors declare that there are no financial, competing or conflicts of interest.